Celine has a fish tank in the shape of a cube. If the volume of her fish tank is 1,000 cubic inches, what is the area of one of the sides of Celine's fish tank?
Answer by ankor@dixie-net.com(15656)

Celine has a fish tank in the shape of a cube.
If the volume of her fish tank is 1,000 cubic inches, what is the area of one of the sides of Celine's fish tank?
:
Let x = 1 side of the cube
we know
x^3 = 1000
therefore
x = 10
and
x^2 = area of one side
10^2 = 100 sq/in